40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14871     Entry time: Wed Sep 11 10:26:56 2019     In reply to: 14868     Reply to this: 14872   14876  
Author: aaron 
Type: Update 
Category: IOO 
Subject: WFS measurements 

Gameplan

We should also have a plan for the next couple weeks so we are organized; heavily adapted from. Here's what I'm thinking this morning:

  1. Construct the input/output matrix for the WFS. (basically, what we did yesterday)
    1. Measure a transfer function of MC[1, 2, 3]_[PIT, YAW] to [WFS1, WFS2, MC2_TRANS]_[PIT, YAW]. The transfer function above the loop bandwidth (few seconds BW, so we will excite >~ 10 Hz) characterizes the response of the sensor to the excitation.
    2. Invert the resulting 3x3 matrix and populate the inverted matrix at WFS_OUTMATRIX. This will map the WFS basis to the MC optics' pit/yaw basis.
    3. Script this process. If we make changes (for example, moving the telescoping lenses) to make this matrix more diagonal, we'll want to do these steps many times.
  2. Characterizing the loop
    1. Optimize the demodulation phase -- we want to minimize the signal in Q. This should also be automated. I found documentation in the white Wave Front Sensing binder
      1. Misalign a mirror in pitch or yaw, and rotate the phase to minimize the magnitude of Q (maximize I); this angle is 'R' on the WFSx_SETTINGS screen.
    2. We should measure a step response applied to each angular dof of the MC optics.
    3. Guoy Phase Calibration
  3. Characterizing / Calibrating the WFS heads
    1. The DCC has LIGO test procedures for their WFS RFPD, as does the white binder; the following checks are relevant for our WFS, and this is how I think we should carry them out (not identical to the procedure as written in the document). For many of these, we'll want to set up the JenneAM laser with a network analyzer for RF modulation.
      1. DC path transimpedance
        1. Measure the DC power of JenneAM with a power meter, and direct the beam to each of the QPD quadrants. Make sure the beam fits on a single quadrant.
        2. This will give us the product of the PD efficiency and DC transimpedance gain
        3. Last time this was measured (white WFS binder)
      2. notch tuning -- we are going to measure the TF, but I won't tune it without someone as ancient as the electronics
        1. Using the network analyzer, measure a transfer function from the laser AM to the QPD head's RF output
          1. Is there a pickoff available? The LIGO testing procedures recommend a FET probe
          2. We should do this while measuring the DC transimpedance for each quadrant
      3. notch rejection ratios
        1. While taking the RF transfer function, use the delta marker to record the difference between the notch and the RF operating frequency.
      4. RF transimpedance
        1. Illuminate the PD with white light from an incandescent bulb (a shot-noise limited source)
          1. 6-10 mA of photocurrent should be generated
        2. Use an RF spectrum analyzer and low noise RF pre-amplifier (gain ~20dB) to measure the shot noise limited spectrum
        3. A piece of scotch tape can be used to make the light uniformly illuminate the QPD
        4. Convert this RF PSD to an rms amplitude (voltage) spectral density, and also note the DC photocurrent. This can be used to calculate the RF transimpedance with
          1. Z_\mathrm{RF} = \sqrt{\frac{V_\mathrm{rms}^2I_\mathrm{DC}}{3.2\times 10^{-19}}}
      5. Shot noise limited input sensitivity
        1. Measure the RF PSD with the beam blocked and light off; this is the dark photocurrent, and can be used to calculate the shot noise limited sensitivity.

References:

  • Binders of documents about the 40m WFS
  • LIGO ISC WFS RFPD test procedure (T1200347 is dual frequency, T1200380 is single frequency)
    • The associated datasheet template is in T1200381
  • Wavefront Sensor (T960111). This document even has a calibration protocol with forms to fill in during testing, so I've printed an extra copy of that appendix.

Automation

It would be good to script some of what we did yesterday. I'm checking out some scripts I'd used for Qryo and armloss measurements to remember the best way to do this.

  • Existing WFS scripts (I didn't try these)
    • WFS_DC_offsets -- sets the WFS QPD dark offsets
      • block beam, then run script
    • MC2_TRANS_offsets -- sets the MC2 transmission offset (why isn't this in the same script as WFS_DC_offsets?)
      • MC should be aligned, beams centered on WFS, WFS servo off
    • mcWFSallowOn(Off) -- turns on (off) the ASC filter module outputs
    • mcwfshold -- turns off the input to WFS servos, but holds the current values of MC optic biases
    • mcwfsoff -- turns off the mc wfs loop
      • First, turns off the WFS outputs (eg WFS1_PIT OUTPUT)
      • Turns off the MC WFS input gains
      • Holds the WFS loop outputs

Miscellany

I noticed yesterday that the PSL_shutterqst box is white, and I've seen timeout requests when eg the reboot script tries to open/close the PSL shutter. It seems like a shutter that should open, so I should find the aux machine to restart it.

ELOG V3.1.3-